# Random-sketching Techniques to Enhance the Numerically Stability of Block Orthogonalization Algorithms for $s$-step GMRES

### v1

@@ 7.2 Randomized BCGS
Flop Count Projection Normalization Total One-stage: standard $2nm(m+1/2)$ $2nm(m+1/2)$ $3nm$ $3nm$ $2nm^{2}$ $2nm^{2}$ $s$-step $2nm(s+1)(m-s+1/2)/s$ $2nm(s+1)(m-s+1/2)/s$ $4nm(s+1)(s+3/2)/s$ $2nm(s+1)(s+3/2)/s$ $2nm^{2}(s+1)/s$ $2nm^{2}(s+1)/s$ sketch($\hat{s}=s$) $2nm(s+1)(m-s+1/2)/s$ $2nm(s+1)(m-s+1/2)/s$ $4nm(s+1)(7s/4+2)/s$ $2nm(s+1)(s+3/2)/s$ $2nm^{2}(s+1)/s$ $2nm^{2}(s+1)/s$ Two-stage: sketch($s<\hat{s}<m$) $2nm(s+1)(m-\hat{s}+1/2)/s$ $2nm(\hat{s}+1)(m-\hat{s}-1/2)/\hat{s}$ $2nm(s+1)(5\hat{s}/2-s/2+5/2)/s$ $2nm(\hat{s}+1)(\hat{s}+3/2)/\hat{s}$ $2nm^{2}(s+1)/s$ $2nm^{2}(\hat{s}+1)/\hat{s}$ $+2nm(\hat{s}+1)(\hat{s}+3/2)/\hat{s}$ sketch($\hat{s}=m$) $2nm(s+1)(5m/2-s/2+5/2)/s$ $nm(s+1)^{2}/s$ $2n(m+1)(m+3/2)$ $5nm^{2}(s+1)/s$ $2nm^{2}$
@@
Table 2: Asymptotic computational complexity of orthogonalization scheme over one GMRES restart cycle, where for a given starting unit vector, the “standard” and “$s$-step” are based on CGS2 and BCGS2 with CholQR2, respectively. For the complexity with the random sketch, we use Gaussian Sketch with the sketch size of $\widehat{m}=2(\widehat{s}+1)$. For each components, there are two columns for each component, “Projection”, “Normalization”, and “Total”, where the first and second columns are for orthogonalization and then for reorthogonalization for the standard and $s$-step algorithms, while they are for the preprocessing based on Gaussian random-sketching and then for orthogonalization for the remaining algorithms. For two-stage, “Projection” and “Normalization” are the inter-block and intra-block orthogonalization of big panels. “Total” only shows the leading terms of the complexity.
@@ 7.3 Remarks on Complexity
We discuss the complexity of the overall algorithms in Section 8, but we provide a few remarks, comparing the one-stage and two-stage algorithms, here.
Compared to the one-stage algorithm with RandCholQR, the two-stage algorithm with RandBCGS reduces the communication cost by delaying the orthogonalization until $\widehat{s}+1$ basis vectors are generated. However, this two-stage algorithm needs to sketch the big panel with the total of $\widehat{s}+1$ basis vectors, hence requiring a larger sketch size than the one-stage algorithm. For instance, with the Gaussian sketch, its sketch size for the two-stage algorithm is proportional to the number of columns in the big panel, $\widehat{s}+1$, while the one-stage algorithm sketches each panel of $s+1$ basis vectors, and its sketch size is proportional to the panel size, $s+1$. Hence, in the randomized two-stage algorithm, there is a trade-off between the reduced communication cost for the orthogonalization and the increased computational cost for sketching, which we discuss in the next section. [^2] As Count-Gaussian sketching has $O(n)$ complexity, it has the potential to remove this overhead of the randomized two-stage algorithm.
When $\widehat{s}<m$, the two-stage algorithm with these two pre-processing schemes in Section 7 requires about the same computational cost as the one-stage algorithm. On the other hand, when $\widehat{s}=m$, the two-stage algorithm has a lower computational cost than the one-stage algorithm (see the asymptotic complexity discussion in Section 8). This is because if the preprocesing step can keep the well-conditioning of the basis vectors over the whole restart-cycle of $s$-step GMRES, then the reorthogonalization (Line 19) is not needed, reducing the total computational cost of the orthogonalization. In this case, it is also possible to skip the $\ell_{2}$-orthogonalization (i.e.,, CholQR on Line 15), and solve the least-square problem in the sketched space. Nevertheless, in this paper, we will focus on generating the $\ell_{2}$-orthogonal basis vectors, but will show the breakdown of the orthogonalization time, including the CholQR time in Section 11.
@@ 8 Asymptotic Complexity
    The sketching framework performs the sketched transformation to generate the well-conditioned basis vectors, followed by the $\ell_{2}$-orthogonalization of the basis vectors. Hence, the framework replaces the first orthogonalization process of the standard algorithm with the sketched transformation, while the orthogonalization process of the framework performs the same number of flops as the re-orthogonalization process of the standard algorithm.
@@
    With $\widehat{s}=s$ (one-stage), the sketching framework requires additional computation and communication volume but the relative overhead is small (about the order of $s/m$ in the total complexity), which leads to insignificant increase in the orthogonalization time in our performance tests.
@@
    With $\widehat{s}=m$, the sketching framework has the best communication latency cost, but the highest computational overhead, where the total computational cost of the orthogonalization is increased by a factor of 1.75×1.75$\times$1.75 ×. Nevertheless, when the performance is limited by the latency, this computational overhead may not be significant. However, with the Gaussian Sketching, this increases the storage cost by a factor of three, which could limit the use of the dense sketching.
@@
    These storage and computational overheads may be reduced using a smaller sketch size ($s<\widehat{s}<m)$. This reduces the overhead to be $\mathcal{O}(\widehat{s}/m)$, though the reduction in the latency is also reduced.
Overall, in our experiments, we obtained the best performance using $\widehat{s}=m$, unless the solver did not require reorthogonalization on a small number of GPUs.
@@ 9 Numerical Experiments
We compared the orthogonality errors of the proposed block orthogonalization schemes using the default double precision in MATLAB. For these studies, instead of studying the numerical properties of the proposed methods within the $s$-step GMRES, we treat them as a general block orthogonalization scheme and use synthetic matrices as the input vectors. This allows us to control the condition number of the matrix easily. Numerical results showing how the condition numbers of the Krylov vectors could grow, can be found in [6].
We first study how the orthogonality errors grow with the condition numbers of the input vectors for the one-stage algorithms. Figure 9 shows the orthogonality error when CholQR2 or RandCholQR is used for the first intra block orthogonalization of the one-stage BCGS2. Our test matrix is the glued matrix [36] that has the same specified order of the condition number for each panel $V_{j}$ and for the overall matrix $V_{1:12}$. As expected, the orthogonality error of the basis vectors $\widehat{Q}$ was $\mathcal{O}(\epsilon)$ using CholQR2 when the condition number of the input matrix is smaller than $\mathcal{O}(\epsilon)^{-1/2}$. With RandCholQR, the same $\mathcal{O}(\epsilon)$ orthogonality errors were obtained as long as the input matrix is numerically full-rank with the condition number of $\mathcal{O}(\epsilon)^{-1}$, and hence demonstrating superior stability compared to CholQR2.
We now show that the two-stage approach obtains the $\mathcal{O}(\epsilon)$ orthogonality errors when the condition (11) or (14) is satisfied, using BCGS-PIP or RandBCGS, respectively. Figures 10 and 11 show the orthogonality errors using the two-stage approach with BCGS-PIP and RandBCGS as the preprocessing schemes, respectively. The test matrix is the glued matrix, where each big panel has the condition number $\mathcal{O}(10^{15})$. For this synthetic matrix, BCGS-PIP failed when the condition number of the accumulated panels increased more than $\mathcal{O}(\epsilon)^{-1/2}$. In contrast, RandBCGS managed to keep the $\mathcal{O}(1)$ condition number of the big panel $[Q_{1:\ell-1},\widehat{Q}_{\ell:t}]$, and the overall orthogonality error of $Q$ was $\mathcal{O}(\epsilon)$.
(a) Condition number (marker at every $s$ steps).
Figure 10: Condition number and orthogonality error using two-stage approach with BCGS-PIP on glued matrix with $(n,m,\hat{s},s)=(100000,180,60,5)$.
(a) Condition number (marker at every $s$ steps).
(b) Orthogonality errors (orange circle marker at every $s$ steps, while green triangle marker at every $\widehat{s}$ steps).
Figure 11: Condition number and orthogonality error using two-stage approach with rand-BCGS2 on glued matrix with $(n,m,\hat{s},s)=(100000,180,60,5)$.
@@ 10 Implementation
We have implemented the random-sketching using standard linear algebra kernels as discussed in Section 5.2. These distributed-memory kernels are readily available through Tpetra or Kokkos-Kernels, given that the random-sketching matrix $\boldsymbol{\Theta}$ is explicitly generated and stored in memory.
@@ 11 Performance Experiments
We conducted our performance tests on the Perlmutter supercomputer at National Energy Research Scientific Computing (NERSC) Center. Each compute node of Perlmutter has one 64-core AMD EPYC 7763 CPUs and four NVIDIA A100 GPUs. On each node, we launched 4 MPI processes (one MPI per GPU) and assigned 16 CPU cores to each MPI process. For all the performance studies, we used the sketch size of $2\hat{s}$ for the Gaussian Sketch, while the sketch size of $2\hat{s}^{2}$ is used for the Count Sketch.
@@
We now study the performance of the block orthogonalization kernels. Firgure 13 show the breakdown of time needed for CholQR on a single NVIDIA A100 GPU. With our experiment setups, the dense triangular solves (TRSM) required for generating the orthogonal basis vectors took longer than the dense matrix-matrix vector multiply (GEMM) needed to generate the Gram matrix. The time needed to compute the Cholesky factorization of the Gram matrix on the CPU was negligible.
@@
Figure 14(a) shows the breakdown of the intra-block orthogonalization time on the four NVIDIA A100 GPUs available on a single Perlmutter compute node. Compared to the Gaussian Sketch, Count-Sketch was slightly faster, but due to its larger sketch size, it required more time for the global-reduce and local Householder QR. Overall, Count-Gaussian sketch obtained the best performance, but the performance on Perlmutter was largely dominated by the global-reduce and TRSM (and not by the sketching time), and Gaussian and Count-Gaussian sketch obtained similar performance.
@@ 11.3 $s$-step GMRES Strong-scaling Results
Table 4 shows the parallel strong-scaling performance of the $s$-step GMRES for solving a 3D Laplace problem. We used the restart length of 100 (i.e., $m=100$), and considered GMRES to have converged when the relative residual norm is reduced by six orders of magnitude.
@@ 12 Conclusion
We integrated random sketching techniques into the block orthogonalization process, required for the $s$-step GMRES. The resulting algorithm ensures that the orthogonalization errors are bounded by the machine precision as long as each of the block vectors are numerically full-rank. Our performance results demonstrated that the numerical stability of the block orthogonalization process is improved with a relatively small performance overhead. Our implementation of the random sketching utilizes standard linear algebra kernels such that it is portable to different computer architectures. Though the vendor-optimized versions of these kernels are often available, they may not be optimized for the specific shapes or sparsity patterns of the sketching matrices, and we have observed the sparse sketch often obtains the suboptimal performance. Nevertheless, the sparse random sketching has the complexity of $\mathcal{O}(n)$, and with a careful implementation, RandCholQR might not only enhance the numerical stability but also be able to outperform CholQR2.

### v2

@@ 7.2 Randomized BCGS
Flop Count Projection Normalization Total One-stage: standard $2nm(m+1/2)$ $2nm(m+1/2)$ $3nm$ $3nm$ $2nm^{2}$ $2nm^{2}$ $s$-step $2nm(s+1)(m-s+1/2)/s$ $2nm(s+1)(m-s+1/2)/s$ $4nm(s+1)(s+3/2)/s$ $2nm(s+1)(s+3/2)/s$ $2nm^{2}(s+1)/s$ $2nm^{2}(s+1)/s$ sketch($\hat{s}=s$) $2nm(s+1)(m-s+1/2)/s$ $2nm(s+1)(m-s+1/2)/s$ $4nm(s+1)(7s/4+2)/s$ $2nm(s+1)(s+3/2)/s$ $2nm^{2}(s+1)/s$ $2nm^{2}(s+1)/s$ Two-stage: pip $2nm(s+1)(m-s+1/2)/s$ $2nm(s+1)(m-s+1/2)/s$ $4nm(s+1)(s+3/2)/s$ $2nm(s+1)(s+3/2)/s$ $2nm^{2}(s+1)/s$ $2nm^{2}(s+1)/s$ sketch($s<\hat{s}<m$) $2nm(s+1)(m-\hat{s}+1/2)/s$ $2nm(\hat{s}+1)(m-\hat{s}-1/2)/\hat{s}$ $2nm(s+1)(5\hat{s}/2-s/2+5/2)/s$ $2nm(\hat{s}+1)(\hat{s}+3/2)/\hat{s}$ $2nm^{2}(s+1)/s$ $2nm^{2}(\hat{s}+1)/\hat{s}$ $+2nm(\hat{s}+1)(\hat{s}+3/2)/\hat{s}$ sketch($\hat{s}=m$) $2nm(s+1)(5m/2-s/2+5/2)/s$ $nm(s+1)^{2}/s$ $2n(m+1)(m+3/2)$ $5nm^{2}(s+1)/s$ $2nm^{2}$
@@
Table 2: Asymptotic computational complexity of orthogonalization scheme over one GMRES restart cycle, where for a given starting unit vector, the “standard” and “$s$-step” are based on CGS2 and BCGS2 with CholQR2, respectively. For the complexity with the random sketch, we use Gaussian Sketch with the sketch size of $\widehat{m}=2(\widehat{s}+1)$. For each components, there are two columns for each component, “Projection”, “Normalization”, and “Total”, where for the standard and $s$-step algorithms, the first and second columns are for orthogonalization and then for reorthogonalization, while for the remaining algorithms, they are for the preprocessing based on Gaussian random-sketching and then for orthogonalization. For two-stage, “Projection” and “Normalization” are the inter-block and intra-block orthogonalization of big panels. “Total” only shows the leading terms of the complexity.
@@ 7.3 Remarks on Complexity
We discuss the overall complexity of the various block orthogonalization algorithms in Section 8, but we provide a few remarks on the complexity, comparing the one-stage and two-stage algorithms, here.
Compared to the one-stage algorithm with RandHH, the two-stage algorithm with RandBCGS2 reduces the communication cost by delaying the orthogonalization until $\widehat{s}+1$ basis vectors are generated. However, this two-stage algorithm needs to sketch the big panel with the total of $\widehat{s}+1$ basis vectors, hence requiring a larger sketch size than the one-stage algorithm. For instance, with the Gaussian sketch, its sketch size for the two-stage algorithm is proportional to the number of columns in the big panel, $\widehat{s}+1$, while the one-stage algorithm sketches each panel of $s+1$ basis vectors, and its sketch size is proportional to the panel size, $s+1$. Hence, in the randomized two-stage algorithm, there is a trade-off between the reduced communication cost for the orthogonalization and the increased computational cost for sketching, which we discuss in the next section. [^2] As Count-Gaussian sketching has $O(n)$ complexity, it has the potential to remove this overhead of the randomized two-stage algorithm.
When $\widehat{s}<m$, the two-stage algorithm with the two pre-processing schemes in Section 7 requires about the same computational cost as the one-stage algorithm. On the other hand, when $\widehat{s}=m$, the two-stage algorithm has a lower computational cost than the one-stage algorithm (see the asymptotic complexity discussion in Section 8). This is because if the preprocesing step can keep the well-conditioning of the basis vectors over the whole restart-cycle of $s$-step GMRES, then the reorthogonalization (Line 19) is not needed, reducing the total computational cost of the orthogonalization. In this case, it is also possible to skip the $\ell_{2}$-orthogonalization (i.e., CholQR on Line 15), and solve the least-square problem in the sketched space. Nevertheless, in this paper, we will focus on generating the $\ell_{2}$-orthogonal basis vectors, but will show the breakdown of the orthogonalization time, including the CholQR time, in Section 11.
@@ 8 Asymptotic Complexity
    The randomized block orthogonalization algorithm performs the preprocessing based on random-sketching to generate the well-conditioned basis vectors, followed by the $\ell_{2}$-orthogonalization of the basis vectors. Hence, the randomized algorithm replaces the first orthogonalization process of the standard algorithm with the randomized transformation of the basis vectors, while the orthogonalization process is identical and performs the same number of flops as the re-orthogonalization process of the standard algorithm.
@@
    With $\widehat{s}=s$ (one-stage), compared to CholQR2, RandCholQR has higher computation complexity and larger communication volume, but the relative overhead is small (about the order of $s/m$) in the total complexity. This leads to insignificant increase in the orthogonalization time (for improving the numerical stability) in our performance tests.
@@
    One-stage algorithm with CholQR2 and two-stage algorithm with BCGS-PIP have about the same computational complexity.
@@
    With $\widehat{s}=m$, the two-stage framework has the best communication latency cost, but random-sketching has the highest computational overhead, where the total computational cost of the orthogonalization is increased by a factor of 1.75×1.75$\times$1.75 ×. Nevertheless, when the performance is limited by the latency, this computational overhead may not be significant.
    Though the storage cost is increased by a factor of three, which could limit the use of the dense Gaussian sketch, the overhead can be reduced by using sparse Count sketch.
  • 6.
    These storage and computational overheads may be reduced using a smaller sketch size ($s<\widehat{s}<m)$. This reduces the overhead to be $\mathcal{O}(\widehat{s}/m)$, though the reduction in the latency is also reduced. Nevertheless, in our experiments, the two-stage algorithm obtained the best performance using $\widehat{s}=m$.
@@ 9 Numerical Experiments
We compared the orthogonality errors of the proposed block orthogonalization schemes using the default double precision in MATLAB. For these studies, instead of studying the numerical properties of the proposed methods within the $s$-step GMRES, we treat them as general block orthogonalization schemes and use synthetic matrices as the input vectors. This allows us to control the condition number of the matrix easily. Numerical results showing how the condition numbers of the Krylov vectors could grow, can be found in [6].
We first study how the orthogonality errors grow with the condition numbers of the input vectors for the one-stage algorithms. Figure 9 shows the orthogonality error when CholQR2 or RandCholQR are used for the first intra block orthogonalization of the one-stage BCGS2. Our test matrix is the glued matrix [36] that has the same specified order of the condition number for each panel $V_{j}$ and for the overall matrix $V_{1:12}$. As expected, the orthogonality error of the basis vectors $\widehat{Q}$ was $\mathcal{O}(\epsilon)$ using CholQR2 when the condition number of the input matrix is smaller than $\mathcal{O}(\epsilon)^{-1/2}$. With RandCholQR, the same $\mathcal{O}(\epsilon)$ orthogonality errors were obtained as long as the input matrix is numerically full-rank with the condition number of $\mathcal{O}(\epsilon)^{-1}$, and hence demonstrating superior stability compared to CholQR2.
Next, we show that the two-stage approach obtains $\mathcal{O}(\epsilon)$ orthogonality error when the condition (12) or (15) is satisfied, using BCGS-PIP or RandBCGS2, respectively. Figure 10(b) shows the condition number of basis vectors using the two-stage approach with BCGS-PIP or RandBCGS2 as the preprocessing schemes, while Figure 11 shows the orthogonality errors using RandBCGS2. The test matrix is the glued matrix, where each big panel has the condition number $\mathcal{O}(10^{15})$. For this synthetic matrix, BCGS-PIP failed when the condition number of the accumulated panels increased more than $\mathcal{O}(\epsilon)^{-1/2}$. In contrast, RandBCGS2 managed to keep the $\mathcal{O}(1)$ condition number of the big panel $[Q_{1:\ell-1},\widehat{Q}_{\ell:t}]$, and the overall orthogonality error of $Q$ was $\mathcal{O}(\epsilon)$.
(a) BCGS-PIP.
(b) RandBCGS2
Figure 10: Condition number (marker at every $s$ steps) using two-stage approach with BCGS-PIP and RandBCGS2 on glued matrix with $(n,m,\hat{s},s)=(100000,180,60,5)$.
Figure 11: Orthogonality error (orange circle marker at every $s$ steps, while green triangle marker at every $\widehat{s}$ steps) using two-stage approach with RandBCGS2 on glued matrix with $(n,m,\hat{s},s)=(100000,180,60,5)$.
@@ 10 Implementation
We have implemented the random-sketching using standard linear algebra kernels as discussed in Section 5.2. These distributed-memory or on-node kernels are readily available through Tpetra or Kokkos-Kernels, given that the random-sketching matrix $\boldsymbol{\Theta}$ is explicitly generated and stored in memory.
@@ 11 Performance Experiments
We conducted our performance tests on the Perlmutter supercomputer at National Energy Research Scientific Computing (NERSC) Center. Each compute node of Perlmutter has one 64-core AMD EPYC 7763 CPUs and four NVIDIA A100 GPUs. On each node, we launched 4 MPI processes (one MPI per GPU) and assigned 16 CPU cores to each MPI process. For all the performance studies with one-stage or two-stage algorithm, we used the sketch size of $2s$ or $2\hat{s}$ for the Gaussian Sketch, while the sketch size of $2s^{2}$ or $2\hat{s}^{2}$ is used for the Count Sketch, respectively.
@@
We now study the performance of the block orthogonalization kernels. Firgure 13 show the breakdown of time needed for CholQR on a single NVIDIA A100 GPU. With our experiment setups, the dense triangular solves (TRSM) required for generating the orthogonal basis vectors took longer than the dense matrix-matrix vector multiply (GEMM) needed to generate the Gram matrix, even though their computational complexity costs are about the same. The time needed to compute the Cholesky factorization of the Gram matrix on the CPU was negligible.
@@
Figure 14(a) shows the breakdown of the intra-block orthogonalization time on the four NVIDIA A100 GPUs available on a single Perlmutter compute node. Compared to the Gaussian Sketch, the “sketch time” (i.e., the time to apply the sketch) of Count-Sketch was slightly faster, but due to its larger sketch size, it required more time for the global-reduce and local Householder QR. Overall, Count-Gaussian sketch obtained the best performance, but the performance on Perlmutter was largely dominated by the global-reduce and TRSM (and not by the sketching time), and Gaussian and Count-Gaussian sketch obtained similar performance.
@@ 11.3 $s$-step GMRES Strong-scaling Results
Although the breakdown of the iteration time depends on the matrix, in Table 4, we show the parallel strong-scaling performance of the $s$-step GMRES for solving a 3D Laplace problem, from which we can infer the performance for other problems (please see, for instance [6], for the performance using difference matrices). We used the restart length of 100 (i.e., $m=100$), and considered GMRES to have converged when the relative residual norm is reduced by six orders of magnitude.
@@ 12 Conclusion
We integrated random sketching techniques into the block orthogonalization process, required for the $s$-step GMRES. The resulting algorithm ensures that the overall orthogonalization errors are bounded by the machine precision as long as each of the block vectors are numerically full-rank. Our performance results demonstrated that the numerical stability of the block orthogonalization process is improved with a relatively small performance overhead. Our implementation of the random sketching utilizes standard linear algebra kernels such that it is portable to different computer architectures. Though the vendor-optimized versions of these kernels are often available, they may not be optimized for the specific shapes or sparsity patterns of the sketching matrices, and we have observed the sparse sketch often obtains the suboptimal performance. Nevertheless, the sparse random sketching has the complexity of $\mathcal{O}(n)$, and with a careful implementation, RandCholQR, using a spares sketch, might not only enhance the numerical stability but also be able to outperform CholQR2.
Appendix
Here we provide a sketch of the proof of Proposition 7.3.
Proof.
Because of Proposition 7.1, it is sufficient to prove that the condition number of the big panel $\widehat{Q}_{\ell+1:t+1}$, before the CholQR (on Line 7 of Figure 6), is bounded as
$\kappa\left(\widehat{Q}_{\ell+1:t+1}\right)=O(1).$ (16)
Then with the bound (7), after the CholQR, the resulting big panel has the orthogonality error,
$\|I-Q_{\ell+1:t+1}^{T}Q_{\ell+1:t+1}\|=O(\epsilon),$
and by an argument identical to [7, Theorem 6.1], the overall stability of the two-stage BCGS2 is ensured with the $\mathcal{O}(\epsilon)$ orthogonality error among all the generated basis vectors,
$\|I-Q_{1:t+1}^{T}Q_{1:t+1}\|=O(\epsilon),$
up to a constant term $c(\epsilon,n,\widehat{s})$.
The general strategy for proving (16) is based on the observation that RandBCGS2 is identical to RandHH except that BCGS2-HH (Line 5 of Figure 8(b)) is used to orthogonalize the sketched vectors $\mathbf{v}_{1:t+1}$ , instead of HH (Line 4 of Figure 4(a)). For instance, the Lines 8–11 of Figure 8(b) are equivalent to applying the forward-substitution to the big panel $\widehat{Q}_{1:t+1}=V_{1:t+1}R_{1:t+1,1:t+1}^{-1}$. Hence, if we show that the backward error incurred during BCGS2-HH differs from HH by a constant factor $c(\epsilon,n,\widehat{s})$, then we can use the error analysis of RandHH from [33] to prove (16).
More specifically, in [33, Corollary 5.1], it was proven that RandHH results in $\kappa\left(\widehat{Q}_{\ell+1:t+1}\right)=O(1)$. This proof relies on [33, Sections 5.2.1–5.2.8]. Now, the error analysis in the first three subsections [33, Sections 5.2.1–5.2.3] applies to RandBCGS2 because they do not depend on the results of the QR factorization. Furthermore, the analysis of the backward error with the QR factorization in [33, Section 5.2.4] only differs by a constant factor if we prove the backward error of the QR factorization via BCGS2-HH only differs from the Householder QR backward error by a constant factor. Hence, the remaining error analysis in [33, Sections 5.2.5–5.2.8] applies to RandBCGS2, which uses BCGS2-HH in place of Householder QR, up to a constant factor. Therefore, by [33, Corollary 5.1], we have $\kappa\left(\widehat{Q}_{\ell+1:t+1}\right)=O(1)$ with RandBCGS2, thereby completing the proof.
The main component of the proof is to derive the bound
$\|\mathbf{v}_{1:j}-\mathbf{\bar{q}}_{1:j}R_{1:j,1:j}\|_{2}=O(\epsilon)\|\mathbf{v}_{1:j}\|_{2},$
where $\mathbf{\bar{q}}_{1:j}$ is an exactly orthogonal matrix and $R_{1:j,1:j}$ is the upper-triangular matrix computed by BCGS2-HH, since such bound was proved for HH in [33, Section 5.2.4]. Since such result is not directly available in [7], we show how to obtain such a bound below.
According to [7, Theorem 6.1], we first note that our computed matrices $\mathbf{q}_{1:j}$ and $R_{1:j,1:j}$ satisfy
$\mathbf{v}_{1:j}+\Delta\mathbf{v}_{1:j}$ $=\mathbf{q}_{1:j}R_{1:j,1:j},$ (17)
$\|\Delta\mathbf{v}_{1:j}\|_{2}$ $=c_{5}(\epsilon,n,\widehat{s})\|\mathbf{v}_{1:j}\|_{2},$ (18)
$\|I-\mathbf{q}_{1:j}^{T}\mathbf{q}_{1:j}\|_{2}$ $=c_{6}(\epsilon,n,\widehat{s}),$ (19)
for some constants $c_{5}(\epsilon,n,\widehat{s})$ and $c_{6}(\epsilon,n,\widehat{s})$.
Given the singular value decomposition of $\mathbf{q}_{1:j}=U\Sigma V^{T}$, its perturbation $\Delta\mathbf{q}_{1:j}$ from the exactly-orthonormal vectors $\bar{\mathbf{q}}_{1:j}$ (i.e., $\Delta\mathbf{q}_{1:j}=\bar{\mathbf{q}}_{1:j}-\mathbf{q}_{1:j}$) can be bounded as
$\|\Delta\mathbf{q}_{1:j}\|_{2}$ $=$ $\|\bar{\mathbf{q}}_{1:j}-\mathbf{q}_{1:j}\|_{2}$ (20)
$=$ $\|U(I-\Sigma)V^{T}\|_{2}$
$=$ $\|I-\Sigma\|_{2}\leq c_{7}(\epsilon,n,\widehat{s}).$
where the singular values of the exactly-orthonormal vectors $\bar{\mathbf{q}}_{1:j}$ are all ones (i.e., $\bar{\mathbf{q}}_{1:j}=UV^{T}$). Using Weyl’s inequality, (19) implies that for each $i\in\{1,\dots,j\}$,
$\sqrt{1-c_{6}(\epsilon,n,\widehat{s})}\leq\Sigma_{i,i}\leq\sqrt{1+c_{6}(\epsilon,n,\widehat{s})}.$ (21)
Hence, the constant $c_{7}(\epsilon,n,\widehat{s})$ is given by
$c_{7}(\epsilon,n,\widehat{s})=\max\{\sqrt{1+c_{6}(\epsilon,n,\widehat{s})}-1,1-\sqrt{1-c_{6}(\epsilon,n,\widehat{s})}\}=O(\epsilon).$
Now, by substituting $\mathbf{q}_{1:j}=\bar{\mathbf{q}}_{1:j}-\Delta\mathbf{q}_{1:j}$ into (17), we obtain
$\bar{\mathbf{q}}_{1:j}R_{1:j,1:j}=\mathbf{v}_{1:j}+\Delta\mathbf{v}_{1:j}+\Delta\mathbf{q}_{1:j}R_{1:j,1:j},$
which allows us to derive the bound,
$\|R_{1:j,1:j}\|_{2}\leq\frac{1+c_{5}(\epsilon,n,\widehat{s})}{1-c_{7}(\epsilon,n,\widehat{s})}\|\mathbf{v}_{1:j}\|_{2},$
because
$(1-c_{7}(\epsilon,n,\widehat{s}))\|R_{1:j,1:j}\|_{2}$
$\;=(1-c_{7}(\epsilon,n,\widehat{s}))\|\bar{\mathbf{q}}_{1:j}R_{1:j,1:j}\|_{2}$
$\;=\|\mathbf{v}_{1:j}+\Delta\mathbf{v}_{1:j}+\Delta\mathbf{q}_{1:j}R_{1:j,1:j}\|_{2}-c_{7}(\epsilon,n,\widehat{s})\|R_{1:j,1:j}\|_{2}$
$\;\leq\|\mathbf{v}_{1:j}\|_{2}+\|\Delta\mathbf{v}_{1:j}\|_{2}+(\|\Delta\mathbf{q}_{1:j}\|_{2}-c_{7}(\epsilon,n,\widehat{s}))\|R_{1:j,1:j}\|_{2}$
$\;\leq\|\mathbf{v}_{1:j}\|_{2}+\|\Delta\mathbf{v}_{1:j}\|_{2}$ (22)
$\;\leq\|\mathbf{v}_{1:j}\|_{2}+c_{5}(\epsilon,n,\widehat{s})\|\mathbf{v}_{1:j}\|_{2}$ (23)
$\;=(1+c_{5}(\epsilon,n,\widehat{s}))\|\mathbf{v}_{1:j}\|_{2}.$
where the inequalities (22) and (23) follows due to $\|\Delta\mathbf{q}_{1:j}\|_{2}\leq c_{7}(\epsilon,n,\widehat{s})$ and $\|\Delta\mathbf{v}_{1:j}\|_{2}=c_{5}(\epsilon,n,\widehat{s})\|\mathbf{v}_{1:j}\|_{2}$ from (20) and (17), respectively.
Thus, we finally have
$\|\Delta\mathbf{v}_{1:j}+\Delta\mathbf{q}_{1:j}R_{1:j,1:j}\|_{2}$
$\leq\|\Delta\mathbf{v}_{1:j}\|_{2}+\|\Delta\mathbf{q}_{1:j}\|_{2}\|R_{1:j,1:j}\|_{2}$
$\leq\left(c_{5}(\epsilon,n,\widehat{s})+c_{7}(\epsilon,n,\widehat{s})\frac{1+c_{5}(\epsilon,n,\widehat{s}}{1-c_{7}(\epsilon,n,\widehat{s})}\right)\|\mathbf{v}_{1:j}\|_{2}$
$=c_{8}(\epsilon,n,\widehat{s})\|\mathbf{v}_{1:j}\|_{2},$ (24)
where $c_{8}(\epsilon,n,\widehat{s})=O(\epsilon)$.
Therefore, there is an exactly-orthogonal matrix $\bar{\mathbf{q}}_{1:j}$ such that the backward error $\Delta\bar{\mathbf{v}}_{1:j}$ from BCGS2-HH satisfies
$\mathbf{v}_{1:j}+\Delta\bar{\mathbf{v}}_{1:j}=\bar{\mathbf{q}}_{1:j}R_{1:j,1:j},$
[…]
In contrast, if we compute the Householder QR of $\mathbf{v}_{1:j}$, the standard backward error analysis [40, Theorem 19.4] gives
$\mathbf{v}_{1:j}+\Delta\mathbf{\tilde{v}}_{1:j}=\mathbf{\tilde{q}}_{1:j}\tilde{R}_{1:j,1:j},$
[…]
where $\tilde{R}_{1:j,1:j}$ is the upper-triangular matrix computed by the HH factorization in finite precision, $\mathbf{\tilde{q}}_{1:j}$ is exactly-orthogonal, and $c_{9}(\epsilon,n,\widehat{s})=O(\epsilon)$ is a constant. In other words, the backward error from BCGS2-HH and Householder QR are both bounded by some $O(\epsilon)\|\mathbf{v}_{1:j}\|_{2}$ terms, and therefore they only differ by a constant factor $c(\epsilon,n,\widehat{s})$. ∎